# Reasoning Beyond Limits: Advances and Open Problems for LLMs

### v1

@@ IV-J8 Critique-out-Loud Reward Models
Critique-out-Loud (CLoud) reward models [95] introduce a generative approach to reward modeling in RLHF by explicitly generating natural language critiques for each response before assigning a scalar reward. Unlike traditional reward models that rely on a single forward pass to evaluate response quality implicitly, CLoud leverages an LM head to produce detailed critiques and a reward head to predict rewards based on the prompt, response, and generated critique. This method, implemented through supervised fine-tuning on oracle critiques from Llama-3.1-405B-Instruct followed by dataset augmentation with self-generated critiques, yields significant improvements—boosting pairwise preference classification accuracy on RewardBench by 4.65% for 8B models and 5.84% for 70B models, and achieving a Pareto improvement in win rate on ArenaHard when used in Best-of-N scoring. Using self-generated (on-policy) critiques and prompts from UltraFeedback and UltraInteract, CLoud demonstrates a promising enhancement in reward modeling that allows for a more explicit and practical evaluation of model responses.
@@
IV-J9 Generative Reward Models
@@
Zhang et al. [96] introduce GenRM, a generative verifier for large language models that leverages the ubiquitous next-token prediction objective to jointly train on verification and solution generation, as opposed to traditional discriminative classifiers used in Best-of-N methods. By doing so, GenRM seamlessly integrates with instruction tuning and chain-of-thought reasoning and can utilize additional inference compute through majority voting, resulting in a 16–64% improvement in problem-solving accuracy on algorithmic and grade-school math tasks compared to standard approaches. The method trains an LLM on synthetic chain-of-thought data to label each section of a response as ”correct” or ”incorrect”—ultimately outputting an overall ”Yes” or ”No”—and uses a maj@K strategy to average token probabilities for final scoring. Notably, fine-tuned small GenRMs have been shown to outperform larger judge models and performance scales favorably with additional training data (an extra  160,000 examples) and model capacity, offering a promising alternative to conventional reward models akin to OpenAI’s CriticGPT approach.
IV-J10 Meta-Rewarding Language Models
@@
Wu et al. [97] propose a self-improvement approach for large language models that emphasizes enhancing both response generation and the model’s internal judgment capabilities through an unsupervised meta-rewarding mechanism. Traditional methods typically rely on human data or self-rewarding schemes focused solely on improving responses, often leading to early saturation in the model’s ability to evaluate its outputs. In contrast, this approach allows a single LLM to assume three distinct roles—actor, judge, and meta-judge—where the actor generates responses, the judge scores them via multiple evaluations, and the meta-judge assesses the quality of these judgments to create refined preference pairs. By using Direct Preference Optimization (DPO) on these actor and judge preference pairs, the model significantly improves its ability to both generate and evaluate responses, as evidenced by substantial performance gains on challenging benchmarks: the win rate on AlpacaEval2 increased from 22.9% to 39.4% and on Arena-Hard from 20.6% to 29.1%. This meta-rewarding strategy, combined with careful length-control to prevent response length explosion, highlights a promising path for self-supervised, iterative model improvement without costly human feedback.
IV-J11 Best-of-N Distillation (BOND)
@@
Sessa et al. [98] propose a novel reinforcement learning from human feedback (RLHF) algorithm, Best-of-N Distillation (BOND), which aims to replicate the benefits of Best-of-N sampling while eliminating its heavy inference-time computational cost. BOND achieves this by aligning the output distribution of the policy model with that of an ideal Best-of-N distribution using a distribution matching framework. Central to this approach is the use of Jeffreys divergence—a balanced combination of forward and backward KL divergences—to navigate the trade-off between exploring a diverse set of outputs (mode-covering) and honing in on high-quality outputs (mode-seeking). The algorithm employs an iterative process that leverages a moving anchor model as a reference point, updated via exponential moving averages, to guide the policy’s improvement. Empirical evaluations on tasks like abstractive summarization and with models from the Gemma family reveal that BOND, particularly its J-BOND variant as used in fine-tuning Google DeepMind’s Gemma 1.1 2B and 7B models, outperforms traditional RLHF methods such as REINFORCE in terms of the reward/KL trade-off.
IV-J12 Simple Preference Optimization (SimPO)
@@
Meng et al. [99] introduce SimPO (Simple Preference Optimization), a novel RLHF method that refines the Direct Preference Optimization (DPO) framework by streamlining the reward formulation. Instead of relying on a separate reference model, SimPO uses the average log probability of a generated sequence as an implicit, length-normalized reward, and it incorporates a target reward margin into the Bradley-Terry objective to enforce a more significant distinction between winning and losing responses. This design reduces computational and memory overhead—cutting processing time by about 20% and GPU memory usage by roughly 10% compared to traditional DPO—and significantly improves performance. Evaluations across multiple state-of-the-art training setups (including models like Mistral, Llama 3, and Gemma 2) and benchmarks such as AlpacaEval 2, MT-Bench, and Arena-Hard reveal that SimPO outperforms conventional approaches by up to 6.4 points on AlpacaEval 2 and 7.5 points on Arena-Hard. Notably, the top-performing model built on Gemma-2-9B-it achieved a 72.4% length-controlled win rate on AlpacaEval 2, a 59.1% win rate on Arena-Hard, and ranked first on Chatbot Arena among sub-10B models with real user votes.
@@
Figure 2: Thought Preference Optimization [100] – The process begins by instructing the language model to produce an internal reasoning sequence before generating its final answer. After multiple potential responses are sampled, these answers are passed to an evaluator model that identifies the most and least preferred options. The complete outputs corresponding to these top and bottom choices serve as accepted and rejected examples for DPO optimization [115]. This cycle is repeated over multiple training iterations.
IV-J13 Thought Preference Optimization (TPO)
@@
Wu et al. [100] proposed a novel training methodology designed to endow LLMs with explicit ”thinking” capabilities before generating their final responses, as presented in Fig. 2. Traditionally, LLMs are optimized to answer user queries directly, but they often lack a structured, intermediate reasoning process crucial for complex tasks. The proposed approach—termed Thought Preference Optimization (TPO)—leverages an iterative search and optimization procedure that generates multiple chain-of-thought (CoT) sequences, explicitly separating the model’s internal reasoning (”thoughts”) from its final answers. For each instruction, a judge model (e.g., ArmoRM) scores the final responses, and preference pairs are constructed by comparing the highest and lowest-scoring outputs. These pairs are then used in a length-controlled Direct Preference Optimization (DPO) training loop, repeated for several iterations to progressively refine the model’s ability to ”think” before answering. Empirical results demonstrate that this procedure yields substantial improvements, with approximately 20% performance gains on benchmarks such as AlpacaEval and Arena-Hard across reasoning-intensive tasks and domains traditionally relying on more straightforward responses, like marketing, health, and general knowledge. While initial performance may temporarily lag behind direct response generation, the iterative process leads to significant overall gains, and TPO outperforms standard supervised fine-tuning on preferred outputs by around 3–4%. Notably, the reward model evaluates only the final responses, underscoring the importance of a separate, optimized thought process. This design choice may reduce performance on datasets like GSM8K if the training data or judge model is not perfectly aligned.
IV-J14 Proximal Policy Optimization (PPO)
@@
Proximal Policy Optimization (PPO) [24] is a reinforcement learning algorithm that achieves robust policy improvement using a clipped surrogate objective to constrain updates and maintain training stability. Unlike Trust Region Policy Optimization (TRPO), which relies on complex second-order methods, PPO employs a first-order approach that limits the ratio of the new policy to the old policy within a predefined range, thereby preventing drastic changes that could destabilize learning.
@@
Let’s consider a pipeline that consists of three stages: supervised fine-tuning (SFT), preference model training, and reinforcement learning via Proximal Policy Optimization (PPO) [116, 117].
@@
[…]
@@
IV-J15 Group Relative Policy Optimization (GRPO)
DeepSeek team propose DeepSeekMath 7B [49], a model specifically designed to tackle the challenges of mathematical reasoning by building on DeepSeek-Coder-Base-v1.5 7B and leveraging an enormous corpus of 120 billion math-related tokens alongside natural language and code data. Notably, DeepSeekMath 7B achieves an impressive 51.7% accuracy on the MATH benchmark without the need for external toolkits or voting techniques, and its performance further rises to 60.9% with self-consistency over 64 samples, approaching the levels of high-end models like Gemini-Ultra and GPT-4. A key innovation in this work is the introduction of Group Relative Policy Optimization (GRPO), an RLHF method that refines mathematical reasoning by optimizing the memory usage of the conventional PPO algorithm. GRPO distinguishes itself by eliminating the need for a separate value function and directly integrating the KL-divergence term into the loss function, leading to roughly a 5% improvement on benchmarks such as GSM8K and MATH. Utilizing an iterative approach with 144,000 chain-of-thought prompts and training its reward model via the “Math-Shepherd” process enables more efficient and specialized reasoning, setting a new standard for open-source language models in mathematical problem-solving.
SWE-RL [118] tackles the significant challenge of scaling reinforcement learning-based reasoning to the real-world domain of software engineering, where tasks like debugging and issue resolution require not only technical expertise but also nuanced reasoning over complex, evolving codebases. Unlike previous RL approaches focusing on competitive coding or math, SWE-RL leverages extensive open-source software evolution data—including code snapshots, pull requests, and issue contexts—to train LLMs to recover developers’ reasoning processes autonomously. By employing a lightweight, rule-based reward function that compares generated solutions to ground-truth patches, and optimizing using Group Relative Policy Optimization (GRPO), SWE-RL achieves state-of-the-art performance among medium-sized models, with a 41.0% solve rate on the human-verified SWE-Bench.
@@
Figure 3: Direct Preference Optimization (DPO) [115] – DPO aligns models with human preferences without employing a reinforcement learning stage. Traditional approaches for fine-tuning language models with human feedback [13] typically involve training a reward model on a dataset of prompts and preference annotations, then using RL to discover a policy that maximizes the learned reward. By contrast, DPO directly adapts the policy to fulfill these preferences best using a straightforward classification-based objective, avoiding both explicit reward modeling and RL.
IV-J16 Direct Preference Optimization (DPO)
Rafailov et al. [102] propose Direct Preference Optimization (DPO), a streamlined method for aligning large language models with human preferences that reframe the alignment challenge as a classification problem on preference data. By leveraging a mapping between reward functions and optimal policies, DPO bypasses the need for a separate reward model and complex PPO-based reinforcement learning, thus simplifying implementation and reducing computational and hyperparameter tuning requirements as presented in Fig. 3. The approach enhances stability by avoiding the instabilities typically associated with traditional RLHF and has been empirically shown to achieve comparable or superior performance in controlling response attributes—such as sentiment in the generated text—and improving quality in tasks like summarization and single-turn dialogue.
@@
IV-J17 Iterative length-regularized DPO (iLR-DPO)
Liu et al. [103] examine Direct Preference Optimization (DPO) and explore its iterative extension using online preferences derived from a trained reward model. The authors identify a critical limitation in the conventional iterative DPO approach: while response quality improves, it inadvertently increases verbosity. To mitigate this issue, they propose an enhanced method—iterative length-regularized DPO (iLR-DPO)—which incorporates a penalty on response length. Empirical evaluations reveal that iLR-DPO enables a 7B model to achieve performance comparable to GPT-4, without the trade-off of excessive verbosity. Notably, the model attains a 50.5% length-controlled win rate against GPT-4 Preview on AlpacaEval 2.0 and demonstrates strong results across various benchmarks, including MT-Bench, Arena-Hard, and the OpenLLM Leaderboard. These findings underscore the efficacy of integrating iterative training with length regularization to improve the alignment of language models with human feedback.
@@
IV-J18 Length-Instruction DPO (LIFT-DPO)
@@
Yuan et al. [104] tackle the persistent issue of length bias in aligned instruction-following models. While such models generally outperform unaligned ones, they tend to generate unnecessarily lengthy responses—a consequence of both evaluation biases and training practices. The authors propose that incorporating explicit length constraints into prompts can resolve this ambiguity, thereby improving both the quality and efficiency of generated outputs. To achieve this, they introduce Length-Instruction Fine-Tuning (LIFT), a method that augments conventional instruction-following datasets with explicit length limits and constructs new preference pairs that penalize verbosity. They also establish new benchmarks—AlpacaEval-LI and MT-Bench-LI—to rigorously assess how well models adhere to these length instructions. Their experiments demonstrate that state-of-the-art models like GPT-4 Turbo and Claude 3 often violate length constraints, whereas models fine-tuned with LIFT exhibit substantially lower violation rates and improved win rates in length-controlled evaluations, all without degrading performance on standard tasks. This work not only offers a promising strategy to mitigate length bias but also provides a more controllable approach to generating concise responses in practical applications.
IV-J19 Identity Preference Optimisation (IPO)
@@
Azar et al. [119] critically examine the two key approximations underlying current RLHF methods for aligning large language models with human preferences. Traditionally, these methods assume that pointwise rewards can replace pairwise preference data and that a reward model trained on such rewards can generalize to unseen data. While Direct Preference Optimization (DPO) has been introduced to bypass the second approximation by learning a policy directly from collected data without an explicit reward model, it still depends on the first assumption. To address this, the authors derive a new general objective, IPO, formulated entirely in pairwise preferences, thereby eliminating both approximations. They further explore a special IPO case—by setting I to the identity function—developing an efficient optimization procedure with proven performance guarantees. Empirical results on illustrative examples indicate that this approach not only provides deeper theoretical insights into the behavior of RLHF and DPO but also achieves superior performance compared to DPO, marking a promising direction for more reliable and theoretically grounded preference-based learning.
IV-J20 Kahneman-Tversky Optimisation (KTO)
@@
Ethayarajh et al. [120] explore how alignment objectives for LLMs, such as those used in Direct Preference Optimization (DPO), inherently capture human biases similar to those described in Kahneman and Tversky’s prospect theory—for example, loss aversion. The authors introduce the concept of human-aware losses (HALOs), arguing that these loss functions outperform traditional cross-entropy minimization by implicitly modeling human perception. However, they note that the utility functions assumed by current methods differ from those in prospect theory. To bridge this gap, the authors propose KTO, which directly maximizes the utility of generated outputs based on a Kahneman-Tversky utility model rather than simply optimizing the log-likelihood of preferences. Notably, KTO matches or exceeds the performance of preference-based methods on models ranging from 1B to 30B parameters, even though it learns solely from a binary signal of output desirability.
IV-J21 Decoupled Clip and Dynamic sAmpling Policy Optimization (DAPO)
@@
Yu et al. [101] introduces DAPO—a novel reinforcement learning algorithm designed to enhance the reasoning capabilities of large language models through inference scaling. Unlike previous methods where critical training details were often omitted, DAPO is fully open-sourced, including its code built on the verl framework and a carefully curated dataset (DAPO-Math-17k), which significantly bolsters reproducibility. The approach distinguishes itself through four key innovations: an asymmetric clipping mechanism (Clip-Higher) that prevents entropy collapse; dynamic sampling that filters out prompts yielding trivial outcomes; a token-level policy gradient loss that controls response length while maintaining reasoning integrity; and a length-aware penalty to mitigate reward noise from potentially valid but truncated long responses. These methodological advances enable DAPO, when applied to the Qwen2.5-32B base model, to achieve 50 points on the AIME 2024 benchmark—surpassing competitors like GRPO with 50% fewer training steps—thereby marking a notable step forward in the scalable training of reasoning-focused LLMs.
@@
[…]
@@
IV-K Retrieval-Augmented Generation (RAG)

### v2

@@ VI-J8 Critique-out-Loud Reward Models
Critique-out-Loud (CLoud) reward models [ankner2024critique] introduce a generative approach to reward modeling in RLHF by explicitly generating natural language critiques for each response before assigning a scalar reward. Unlike traditional reward models that rely on a single forward pass to evaluate response quality implicitly, CLoud leverages an LM head to produce detailed critiques and a reward head to predict rewards based on the prompt, response, and generated critique. This method, implemented through supervised fine-tuning on oracle critiques from Llama-3.1-405B-Instruct followed by dataset augmentation with self-generated critiques, yields significant improvements boosting pairwise preference classification accuracy on RewardBench by 4.65% for 8B models and 5.84% for 70B models, and achieving a Pareto improvement in win rate on ArenaHard when used in Best-of-N scoring. Using self-generated (on-policy) critiques and prompts from UltraFeedback and UltraInteract, CLoud demonstrates a promising enhancement in reward modeling that allows for a more explicit and practical evaluation of model responses.
@@
VI-J9 Generative Reward Models
@@
Zhang et al. [zhang2024generative] introduce GenRM, a generative verifier for large language models that leverages the ubiquitous next-token prediction objective to jointly train on verification and solution generation, as opposed to traditional discriminative classifiers used in Best-of-N methods. By doing so, GenRM seamlessly integrates with instruction tuning and chain-of-thought reasoning and can utilize additional inference compute through majority voting, resulting in a 16–64% improvement in problem-solving accuracy on algorithmic and grade-school math tasks compared to standard approaches. The method trains an LLM on synthetic chain-of-thought data to label each section of a response as ”correct” or ”incorrect” ultimately outputting an overall ”Yes” or ”No” and uses a maj@K strategy to average token probabilities for final scoring. Notably, fine-tuned small GenRMs have been shown to outperform larger judge models and performance scales favorably with additional training data (an extra  160,000 examples) and model capacity, offering a promising alternative to conventional reward models akin to OpenAI’s CriticGPT approach.
VI-J10 Meta-Rewarding Language Models
@@
Wu et al. [wu2024meta] propose a self-improvement approach for large language models that emphasizes enhancing both response generation and the model’s internal judgment capabilities through an unsupervised meta-rewarding mechanism. Traditional methods typically rely on human data or self-rewarding schemes focused solely on improving responses, often leading to early saturation in the model’s ability to evaluate its outputs. In contrast, this approach allows a single LLM to assume three distinct roles actor, judge, and meta-judge where the actor generates responses, the judge scores them via multiple evaluations, and the meta-judge assesses the quality of these judgments to create refined preference pairs. By using Direct Preference Optimization (DPO) on these actor and judge preference pairs, the model significantly improves its ability to both generate and evaluate responses, as evidenced by substantial performance gains on challenging benchmarks: the win rate on AlpacaEval2 increased from 22.9% to 39.4% and on Arena-Hard from 20.6% to 29.1%. This meta-rewarding strategy, combined with careful length-control to prevent response length explosion, highlights a promising path for self-supervised, iterative model improvement without costly human feedback.
VI-J11 Best-of-N Distillation (BOND)
@@
Sessa et al. [sessa2024bond] propose a novel reinforcement learning from human feedback (RLHF) algorithm, Best-of-N Distillation (BOND), which aims to replicate the benefits of Best-of-N sampling while eliminating its heavy inference-time computational cost. BOND achieves this by aligning the output distribution of the policy model with that of an ideal Best-of-N distribution using a distribution matching framework. Central to this approach is the use of Jeffreys divergence a balanced combination of forward and backward KL divergences to navigate the trade-off between exploring a diverse set of outputs (mode-covering) and honing in on high-quality outputs (mode-seeking). The algorithm employs an iterative process that leverages a moving anchor model as a reference point, updated via exponential moving averages, to guide the policy’s improvement. Empirical evaluations on tasks like abstractive summarization and with models from the Gemma family reveal that BOND, particularly its J-BOND variant as used in fine-tuning Google DeepMind’s Gemma 1.1 2B and 7B models, outperforms traditional RLHF methods such as REINFORCE in terms of the reward/KL trade-off.
VI-J12 Simple Preference Optimization (SimPO)
@@
Meng et al. [meng2024simpo] introduce SimPO (Simple Preference Optimization), a novel RLHF method that refines the Direct Preference Optimization (DPO) framework by streamlining the reward formulation. Instead of relying on a separate reference model, SimPO uses the average log probability of a generated sequence as an implicit, length-normalized reward, and it incorporates a target reward margin into the Bradley-Terry objective to enforce a more significant distinction between winning and losing responses. This design reduces computational and memory overhead cutting processing time by about 20% and GPU memory usage by roughly 10% compared to traditional DPO and significantly improves performance. Evaluations across multiple state-of-the-art training setups (including models like Mistral, Llama 3, and Gemma 2) and benchmarks such as AlpacaEval 2, MT-Bench, and Arena-Hard reveal that SimPO outperforms conventional approaches by up to 6.4 points on AlpacaEval 2 and 7.5 points on Arena-Hard. Notably, the top-performing model built on Gemma-2-9B-it achieved a 72.4% length-controlled win rate on AlpacaEval 2, a 59.1% win rate on Arena-Hard, and ranked first on Chatbot Arena among sub-10B models with real user votes.
@@
Figure 5: Thought Preference Optimization [wu2024thinking] – The process begins by instructing the language model to produce an internal reasoning sequence before generating its final answer. After multiple potential responses are sampled, these answers are passed to an evaluator model that identifies the most and least preferred options. The complete outputs corresponding to these top and bottom choices serve as accepted and rejected examples for DPO optimization [rafailov2023direct]. This cycle is repeated over multiple training iterations.
VI-J13 Thought Preference Optimization (TPO)
@@
Wu et al. [wu2024thinking] proposed a novel training methodology designed to endow LLMs with explicit ”thinking” capabilities before generating their final responses, as presented in Fig. 5. Traditionally, LLMs are optimized to answer user queries directly, but they often lack a structured, intermediate reasoning process crucial for complex tasks. The proposed approach termed Thought Preference Optimization (TPO) leverages an iterative search and optimization procedure that generates multiple chain-of-thought (CoT) sequences, explicitly separating the model’s internal reasoning (”thoughts”) from its final answers. For each instruction, a judge model (e.g., ArmoRM) scores the final responses, and preference pairs are constructed by comparing the highest and lowest-scoring outputs. These pairs are then used in a length-controlled Direct Preference Optimization (DPO) training loop, repeated for several iterations to progressively refine the model’s ability to ”think” before answering. Empirical results demonstrate that this procedure yields substantial improvements, with approximately 20% performance gains on benchmarks such as AlpacaEval and Arena-Hard across reasoning-intensive tasks and domains traditionally relying on more straightforward responses, like marketing, health, and general knowledge. While initial performance may temporarily lag behind direct response generation, the iterative process leads to significant overall gains, and TPO outperforms standard supervised fine-tuning on preferred outputs by around 3–4%. Notably, the reward model evaluates only the final responses, underscoring the importance of a separate, optimized thought process. This design choice may reduce performance on datasets like GSM8K if the training data or judge model is not perfectly aligned.
VI-J14 Proximal Policy Optimization (PPO)
@@
Proximal Policy Optimization (PPO) [ouyang2022training] is a reinforcement learning algorithm that achieves robust policy improvement using a clipped surrogate objective to constrain updates and maintain training stability. Unlike Trust Region Policy Optimization (TRPO), which relies on complex second-order methods, PPO employs a first-order approach that limits the ratio of the new policy to the old policy within a predefined range, thereby preventing drastic changes that could destabilize learning.
@@
Let’s consider a pipeline that consists of three stages: supervised fine-tuning (SFT), preference model training, and reinforcement learning via Proximal Policy Optimization (PPO) [schulman2017proximal, stiennon2020learning].
@@
$\mathcal{L}_{\mathrm{PPO}}(\phi)$ $=\mathbb{E}\Bigl[\min\Bigl(r_{t}(\phi)\,A_{t},\;\mathrm{clip}\bigl(r_{t}(\phi),\,1-\epsilon,\,1+\epsilon\bigr)\,A_{t}\Bigr)\Bigr],$
@@
VI-J15 Group Relative Policy Optimization (GRPO)
DeepSeek team propose DeepSeekMath 7B [shao2024deepseekmath], a model specifically designed to tackle the challenges of mathematical reasoning by building on DeepSeek-Coder-Base-v1.5 7B and leveraging an enormous corpus of 120 billion math-related tokens alongside natural language and code data. Notably, DeepSeekMath 7B achieves an impressive 51.7% accuracy on the MATH benchmark without the need for external toolkits or voting techniques, and its performance further rises to 60.9% with self-consistency over 64 samples, approaching the levels of high-end models like Gemini-Ultra and GPT-4. A key innovation in this work is the introduction of Group Relative Policy Optimization (GRPO), an RLHF method that refines mathematical reasoning by optimizing the memory usage of the conventional PPO algorithm. GRPO distinguishes itself by eliminating the need for a separate value function and directly integrating the KL-divergence term into the loss function, leading to roughly a 5% improvement on benchmarks such as GSM8K and MATH. Utilizing an iterative approach with 144,000 chain-of-thought prompts and training its reward model via the “Math-Shepherd” process enables more efficient and specialized reasoning, setting a new standard for open-source language models in mathematical problem-solving.
SWE-RL [wei2025swerladvancingllmreasoning] tackles the significant challenge of scaling reinforcement learning-based reasoning to the real-world domain of software engineering, where tasks like debugging and issue resolution require not only technical expertise but also nuanced reasoning over complex, evolving codebases. Unlike previous RL approaches focusing on competitive coding or math, SWE-RL leverages extensive open-source software evolution data including code snapshots, pull requests, and issue contexts to train LLMs to recover developers’ reasoning processes autonomously. By employing a lightweight, rule-based reward function that compares generated solutions to ground-truth patches, and optimizing using Group Relative Policy Optimization (GRPO), SWE-RL achieves state-of-the-art performance among medium-sized models, with a 41.0% solve rate on the human-verified SWE-Bench.
@@
Figure 6: Direct Preference Optimization (DPO) [rafailov2023direct] – DPO aligns models with human preferences without employing a reinforcement learning stage. Traditional approaches for fine-tuning language models with human feedback [christiano2017deep] typically involve training a reward model on a dataset of prompts and preference annotations, then using RL to discover a policy that maximizes the learned reward. By contrast, DPO directly adapts the policy to fulfill these preferences best using a straightforward classification-based objective, avoiding both explicit reward modeling and RL.
VI-J16 Direct Preference Optimization (DPO)
Rafailov et al. [rafailov2024direct] propose Direct Preference Optimization (DPO), a streamlined method for aligning large language models with human preferences that reframe the alignment challenge as a classification problem on preference data. By leveraging a mapping between reward functions and optimal policies, DPO bypasses the need for a separate reward model and complex PPO-based reinforcement learning, thus simplifying implementation and reducing computational and hyperparameter tuning requirements as presented in Fig. 6. The approach enhances stability by avoiding the instabilities typically associated with traditional RLHF and has been empirically shown to achieve comparable or superior performance in controlling response attributes such as sentiment in the generated text and improving quality in tasks like summarization and single-turn dialogue.
@@
VI-J17 Iterative length-regularized DPO (iLR-DPO)
Liu et al. [liu2024iterative] examine Direct Preference Optimization (DPO) and explore its iterative extension using online preferences derived from a trained reward model. The authors identify a critical limitation in the conventional iterative DPO approach: while response quality improves, it inadvertently increases verbosity. To mitigate this issue, they propose an enhanced method iterative length-regularized DPO (iLR-DPO) which incorporates a penalty on response length. Empirical evaluations reveal that iLR-DPO enables a 7B model to achieve performance comparable to GPT-4, without the trade-off of excessive verbosity. Notably, the model attains a 50.5% length-controlled win rate against GPT-4 Preview on AlpacaEval 2.0 and demonstrates strong results across various benchmarks, including MT-Bench, Arena-Hard, and the OpenLLM Leaderboard. These findings underscore the efficacy of integrating iterative training with length regularization to improve the alignment of language models with human feedback.
@@
VI-J18 Length-Instruction DPO (LIFT-DPO)
@@
Yuan et al. [yuan2024following] tackle the persistent issue of length bias in aligned instruction-following models. While such models generally outperform unaligned ones, they tend to generate unnecessarily lengthy responses a consequence of both evaluation biases and training practices. The authors propose that incorporating explicit length constraints into prompts can resolve this ambiguity, thereby improving both the quality and efficiency of generated outputs. To achieve this, they introduce Length-Instruction Fine-Tuning (LIFT), a method that augments conventional instruction-following datasets with explicit length limits and constructs new preference pairs that penalize verbosity. They also establish new benchmarks AlpacaEval-LI and MT-Bench-LI to rigorously assess how well models adhere to these length instructions. Their experiments demonstrate that state-of-the-art models like GPT-4 Turbo and Claude 3 often violate length constraints, whereas models fine-tuned with LIFT exhibit substantially lower violation rates and improved win rates in length-controlled evaluations, all without degrading performance on standard tasks. This work not only offers a promising strategy to mitigate length bias but also provides a more controllable approach to generating concise responses in practical applications.
VI-J19 Identity Preference Optimisation (IPO)
@@
Azar et al. [azar2024general] critically examine the two key approximations underlying current RLHF methods for aligning large language models with human preferences. Traditionally, these methods assume that pointwise rewards can replace pairwise preference data and that a reward model trained on such rewards can generalize to unseen data. While Direct Preference Optimization (DPO) has been introduced to bypass the second approximation by learning a policy directly from collected data without an explicit reward model, it still depends on the first assumption. To address this, the authors derive a new general objective, IPO, formulated entirely in pairwise preferences, thereby eliminating both approximations. They further explore a special IPO case by setting I to the identity function developing an efficient optimization procedure with proven performance guarantees. Empirical results on illustrative examples indicate that this approach not only provides deeper theoretical insights into the behavior of RLHF and DPO but also achieves superior performance compared to DPO, marking a promising direction for more reliable and theoretically grounded preference-based learning.
VI-J20 Kahneman-Tversky Optimisation (KTO)
@@
Ethayarajh et al. [ethayarajh2024kto] explore how alignment objectives for LLMs, such as those used in Direct Preference Optimization (DPO), inherently capture human biases similar to those described in Kahneman and Tversky’s prospect theory for example, loss aversion. The authors introduce the concept of human-aware losses (HALOs), arguing that these loss functions outperform traditional cross-entropy minimization by implicitly modeling human perception. However, they note that the utility functions assumed by current methods differ from those in prospect theory. To bridge this gap, the authors propose KTO, which directly maximizes the utility of generated outputs based on a Kahneman-Tversky utility model rather than simply optimizing the log-likelihood of preferences. Notably, KTO matches or exceeds the performance of preference-based methods on models ranging from 1B to 30B parameters, even though it learns solely from a binary signal of output desirability.
VI-J21 Decoupled Clip and Dynamic sAmpling Policy Optimization (DAPO)
@@
Yu et al. [yu2025dapoopensourcellmreinforcement] introduces DAPO a novel reinforcement learning algorithm designed to enhance the reasoning capabilities of large language models through inference scaling. Unlike previous methods where critical training details were often omitted, DAPO is fully open-sourced, including its code built on the verl framework and a carefully curated dataset (DAPO-Math-17k), which significantly bolsters reproducibility. The approach distinguishes itself through four key innovations: an asymmetric clipping mechanism (Clip-Higher) that prevents entropy collapse; dynamic sampling that filters out prompts yielding trivial outcomes; a token-level policy gradient loss that controls response length while maintaining reasoning integrity; and a length-aware penalty to mitigate reward noise from potentially valid but truncated long responses. These methodological advances enable DAPO, when applied to the Qwen2.5-32B base model, to achieve 50 points on the AIME 2024 benchmark surpassing competitors like GRPO with 50% fewer training steps thereby marking a notable step forward in the scalable training of reasoning-focused LLMs.
@@
$\epsilon_{\mathrm{low}},\;1+\epsilon_{\mathrm{high}}\bigr]\,A_{i}^{j}\Bigr)\biggr],$
@@
VI-K Retrieval-Augmented Generation (RAG)